Please help solve (s) on the following equation (144/s)+2=144/(s-6)
Answer by lwsshak3(11628)

Please help solve (s) on the following equation
(144/s)+2=144/(s-6)
LCD: s(s-6)
(144*(s-6)+2*s(s-6)=144s
144s-864+2s^2-12s=144s
2s^2-12s-864=0
s^2-6s-432=0
(s+18)(s-24)=0
s=-18
or
s=24